OpenStudy (anonymous):

I just want to know if my answer right?? Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 10x – y = 12.5 3x + y = 7

OpenStudy (anonymous):

Eq1: 10X - Y = 12.5. Eq2: 3X + Y = 7. Add the 2 Eqs and get: 13X + 0 = 19.5. X = 1.5. In Eq1, substitute 1.5 for X: 10*1.5 - Y = 12.5, 15 - Y = 12.5, Y = 2.5. Solution set: (X,Y) = (1.5,2.5).

OpenStudy (anonymous):

Just eleminate the y and add the x and the 12.5 and 7

OpenStudy (anonymous):

then after you do that, just divide.

OpenStudy (lalaly):

Thats right Donnice

OpenStudy (anonymous):

ty